OpenStudy (anonymous):

simplify 3 radical 128

OpenStudy (radar):

\[3\sqrt{128}\]\[3\sqrt{64}\sqrt{2}=24\sqrt{2}\] Did you follow that with understanding?

OpenStudy (anonymous):

did you do the greates common factor forc 128

OpenStudy (radar):

No, I looked for the largest factor that was a perfect square, and 64 was it, 128=2*64

OpenStudy (anonymous):

oh thanks

OpenStudy (radar):

You're welcome.